Index Demo

markus schnalke 2014-05-05

Inhaltsverzeichnis

1 Unix 3 Index 3

1 Unix

Unix is an operating system. It was invented in the late 60s and early 70s at the AT&T Bell Labs.

Index

{AT&T, 3 \atop Unix, 3}